Amount of an investment. The amount of an investment
of P dollars for t years at simple interest rate r is given by
A = P + Prt.
a) Rewrite this formula by factoring out the greatest
common factor on the right-hand side.
b) Find A if $8300 is invested for 3 years at a simple
interest rate of 15%.

Answer by stanbon(75887)

A = P + Prt.
a) Rewrite this formula by factoring out the greatest
common factor on the right-hand side.
A = P(1+rt)
----------------
b) Find A if $8300 is invested for 3 years at a simple
interest rate of 15%.
A = 8300(1 + 0.15*3)
A = $12035.00
================
Cheers,
Stan H.